OpenStudy (anonymous):

How would I find the intersection points for this system of equations? (I'll put the equations in the comments)

OpenStudy (anonymous):

\[\frac{ x ^{2} }{ 4 }+y ^{2}=1\]\[y=x+1\]

OpenStudy (anonymous):

Since y = x + 1... then can I plug it into the other equation? \[\frac{ x ^{2} }{ 4 }+(x+1)^{2}=1\]

OpenStudy (anonymous):

thats what i was thinking

OpenStudy (anonymous):

I'm just not sure how to solve for x from there /:

OpenStudy (anonymous):

multiply out the brackets, times through by 4 and solve the equation, im trying to do it also

OpenStudy (anonymous):

iv got my answer -8/5

OpenStudy (anonymous):

5x^2 +8x+4=4 5x^2 + 8x=0 5x + 8 =0 5x = -8 x = -8/5

OpenStudy (anonymous):

x=-8/5 y=-3/5

OpenStudy (anonymous):

I'm not quite sure how you got that... when I mutiplied out the parenthesis I got \[\frac{ x ^{2} }{ 4 }+x^{2}+2x+1=1\] and I moved the 4 to the other side of the equation to get \[2x ^{2}+2x+1=4\]

OpenStudy (anonymous):

But I don't know if I'm doing it right either...

OpenStudy (anonymous):

you have to times through the whole equation by 4 so x^2/4 + x^2+2x+1=1 times everything my 4 so it becomes x^2 +4x^2 + 8x + 4 = 4

OpenStudy (anonymous):

Ohhh I understand now

OpenStudy (anonymous):

And how did you turn 5x^2 + 8x = 0 into 5x + 8 = 0?

OpenStudy (anonymous):

divide by x

OpenStudy (anonymous):

Oh okay

OpenStudy (anonymous):

And how did you find y? Did you plug the x value into the y = x + 1 equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

And how do I find the other intersection? There are two.

OpenStudy (anonymous):

since we only got 1 answer from finding x there is only one intersection

OpenStudy (anonymous):

otherwise we would have gotten a polynomial answer not a linear

OpenStudy (anonymous):

I graphed the equations on a calculator though, and it shows two intersections. The first equation is a circle and the other one is a line.

OpenStudy (anonymous):

x=0 is also an answer, as we divided through by x

OpenStudy (anonymous):

When did we get that? I mean, it's right, but when did we get that?

OpenStudy (anonymous):

dividing through by x also means that x=0